4. Solve the following initial value problems and sketch the solutions: (a) 4y – 4y + y = 0, y(1) = -4, y(1) = 0 (b) y +
Answer #1

4. 44 - 46 ty co solve the IVP g (15r4 . j (1950. det roema Cotol be atrial Satu Then the c auxiliary equation 4 m² - 4m+1=:. The required y (1) = salution is ent 2 e 2x ② 91092-1/2 14/101=52 -282 y + 3y =0 , Then the auriliony reanis mr_ aram +3=